# MiniPLM: Knowledge Distillation for Pre-Training Language Models

### v2

@@ 3.1 Experimental Setup
We adopt the Qwen-1.5 [3] architecture in our experiments. We use the officially released 1.8B Qwen-1.5 model as the teacher LM and distill its knowledge into students with 200M, 500M, and 1.2B parameters. Detailed model configurations are provided in Appendix B.
@@
We construct pre-training corpora from the Pile [23]. To control the computation in experiments, we pre-train all LMs on a maximum of 50B tokens, where documents are merged to construct instances with sequence lengths of 1,024. For online KD methods that incur additional train-time computation, we reduce their training steps to align the total training computation with pre-training without KD or offline KD methods. See Appendix B for more pre-training details.
@@
  • Vanilla KD [58] minimizes the token-level forward KLD between $p$ and $q_{\bm{\theta}}$, which requires online inference of the teacher LM to obtain the token-level output distributions.
  • SeqKD [41] trains the student LM on the teacher-generated data. Since it is infeasible to generate all 50B tokens, we approximate Kim and Rush [41] by using the first 768 tokens of each instance from the training corpus in Pre-Train w/o KD as the prompts and let the teacher LM generate the remaining tokens offline.
  • MiniLLM [32] minimizes the reverse KLD between $p$ and $q_{\bm{\theta}}$ with PPO [68], which requires online inference of the teacher LM and online sampling from the student LM. We treat the first 768 tokens of the instances from the training corpus of Pre-Train w/o KD as the prompts and sample 256 tokens from $q_{\bm{\theta}}$ during the exploration of PPO.
@@
We employ a 104M reference LM trained on 5B tokens. In Section 3.2 and 3.3, we consider a setting where $\mathcal{D}$ is sufficiently large, containing 105B tokens uniformly sampled from the Pile corpus. We reserve 5B tokens as $\mathcal{D}_{\text{ref}}$, and conduct Difference Sampling as per Eq. (4) on the other 100B tokens by setting $\alpha=0.5$ to construct a 50B-token corpus $\mathcal{D}^{\prime}$. In this way, the student LM is pre-trained on $\mathcal{D}^{\prime}$ for one epoch. We use the loss difference of the teacher and reference LM to sampled instances from $\mathcal{D}$, which is equivalent to Eq. (4) as all instances have 1,024 tokens. In Section 3.4, we evaluate MiniPLM in a data-limited setting, where $\mathcal{D}$ is controlled to contain 50B tokens and the student LM is trained on the difference-sampled data for multiple epochs. See Appendix C.4 for the ablation studies on the reference LM and sampling ratio $\alpha$.
@@
We assess the zero-shot accuracy of LMs trained with different methods on 9 downstream tasks widely used in examining the foundation abilities of base models [78, 26]. We also test the language modeling capability of the LMs on a subset of DCLM [45], a high-quality corpus carefully curated with complex pipelines, to examine how well LMs capture broad and diverse knowledge. See Appendix B for more evaluation details.
@@ 1.8B Teacher $\rightarrow$ 500M Student
Table 1: Zero-shot accuracy scores on 9 widely-used downstream tasks and the average scores (Avg.). We use the Qwen-1.5 1.8B LM [3] as the teacher and Qwen LMs with 200M, 500M, and 1.2B parameters as the student. Student LMs with the same sizes consume the same training-time computation. The best scores of each model size are boldfaced.
@@ 3.2 Main Results
Figure 4: Language modeling loss on the DCLM [45] subset. We distill the knowledge of the 1.8B Qwen model [3] into student LMs from the Qwen family with 200M, 500M, and 1.2B parameters. We control the total training-time FLOPs of different methods to be the same.
@@
Table 2: Test loss predictions using the Scaling Law [37]. $N_{\text{stu}}$: the student LM size. $L_{\text{1T}}$, $L_{\text{10T}}$: the loss when Pre-Train w/o KD and MiniPLM process 1T and 10T tokens, with Vanilla KD consuming the same training FLOPs.
@@
Figure 4, compares the language modeling performance of the MiniPLM-trained LMs and baselines on the DCLM [45] subset, a diverse and high-quality dataset curated from web corpora. The results show that, given the same training-time FLOPs, LMs trained with MiniPLM achieve the lowest test losses. In Table 2, we extrapolate the test losses with the Scaling Law [37] to simulate pre-training on 1T and 10T tokens (details in Appendix C.2), showing thatMiniPLM maintains its advantages at the scales of pre-training recent large LMs [78, 19]. A critical stage in DCLM’s data-cleaning pipeline involves removing easy and common patterns, thereby enhancing challenging and diverse signals. Therefore, a lower test loss on DCLM suggests that, with the teacher LM’s guidance and the reference LM, the MiniPLM-trained LMs learn the diverse and hard knowledge better due to the up-sampling of the corresponding parts in the pre-training distribution, as illustrated in Figure 3(b).
@@ 3.3 KD Across Model Families
A noticeable advantage of MiniPLM over Vanilla KD and MiniLLM is its flexibility to distill the knowledge of a teacher LM into student LMs with completely different tokenizers and architectures without additional strategies like Boizard et al. [7]. In Table 5, we illustrate the performance when using the LMs from the Qwen [3] family as the teacher and the reference LM to distill knowledge into a 212M Llama3.1 [19] model and a 140M Mamba [27] model. The results demonstrate the promising performance of MiniPLM in KD across model families, outperforming Pre-Train w/o KD and the existing offline KD baseline (SeqKD). This allows emerging LMs with novel architectures [47, 73] or advanced tokenization [74, 25] to inherit knowledge from existing LMs, thereby facilitating the development of more efficient and higher-performed models.
@@ 3.4 Data-Limited Setting
We evaluate MiniPLM in a data-limited setting where $\mathcal{D}$ is constrained to contain 50B tokens. To this end, the student LM should be trained on the difference-sampled $\mathcal{D}^{\prime}$ over multiple epochs to ensure the total computation and trained tokens remain consistent with Pre-Train without KD. We split a $|\mathcal{D}_{\text{ref}}|$ containing 1B tokens to train the reference LM. Therefore, for a sampling ratio $\alpha$, the student LM should be trained for $\frac{|\mathcal{D}|}{\alpha|\mathcal{D}-\mathcal{D}_{\text{ref}}|}\approx\frac{1}{\alpha}$ epochs, given that $|\mathcal{D}_{\text{ref}}|\ll|\mathcal{D}|$. In Figure 5, we plot the loss curves of the 200M student LMs on the DCLM corpus when using $\alpha\in\left[0.5,0.25,0.125\right]$ and training the LM for around 2, 4, and 8 epochs, respectively. We can see that difference-sampling 25% data (with $\alpha=0.25$) and training the student LM for 4 epochs yields the best performance, which aligns with the observations in Muennighoff et al. [57]. The corpus sampled with a higher $\alpha$ does not achieve the best quality and diversity offered by Difference Sampling, while a lower $\alpha$ leads to rapid over-fitting of the student LM. These findings suggest that MiniPLM is a promising approach to enhance data utilization when high-quality web corpora become scarce [80]. By extrapolating the loss curve of Pre-Train w/o KD using the Scaling Law in a data-constrained setting [57], we estimate that it would require an additional 68B training tokens to match the performance of MiniPLM ($\alpha=0.25$, 4 epochs), which means MiniPLM reduces the pre-training data requirement by 2.4 times. See Appendix C.2 for more details on this extrapolation.
@@ 3.5 Analysis
Impact of Teacher Model.
Intuitively, larger teacher LMs will lead to better KD results, which is observed in recent works [32]. However, early works have also shown that an excessively large gap between teacher and student models can hinder effective KD [55]. In Figure 6, we plot the performance of Vanilla KD and MiniPLM when distilling teacher LMs with different sizes into a 200M student model. We observe a similar phenomenon to Mirzadeh et al. [55] on Vanilla KD and MiniPLM that larger teacher LMs are not necessarily more helpful for KD. However, we attribute this to different factors for Vanilla KD and MiniPLM. In Vanilla KD, the overhead introduced by larger LMs during training diminishes the benefits of distillation. As for MiniPLM, a 500M teacher LM proves most effective for distilling into a 200M student LM. Smaller teacher LMs (e.g., 300M) lack the capacity to identify the hard but valuable parts of the pre-training distribution, weakening the effectiveness of Difference Sampling as shown in Figure 3(b). On the other hand, the value scale of $\log p({\bm{x}})$ from oversized LMs (e.g., 4B) becomes too small compared to that of the reference LM, which tends to degenerate Difference Sampling into sampling with $p_{\text{ref}}({\bm{x}})$ only, losing the effect of the teacher LM. Future research could focus on optimizing teacher model size for MiniPLM or mitigating the impact of differing $\log p({\bm{x}})$ and $\log p_{\text{ref}}({\bm{x}})$ value scales.
@@
To further verify the effectiveness of Difference Sampling on improving the diversity of the pre-training corpus, in Table 6, we follow Friedman and Dieng [22] to compute the semantic diversity of the difference-sampled corpus and the data used in other baseline pre-training approaches. The results show that the difference-sampled corpus has the highest diversity, despite that Difference Sampling is derived from minimizing the reverse KLD, which exhibits the mode-seeking behavior [54]. We suspect the reason is that Difference Sampling down-samples the easy parts of the corpus containing repeated contents while up-sampling the hard parts consisting of diverse texts. These two components, with large $p({\bm{x}})$ values as seen in Figure 3(b), constitute the major modes of the teacher LM that $q_{\bm{\theta}}$ seeks during optimizing Eq. (1). Therefore, the mode-seeking behavior helps remove the noisy parts of the pre-training distribution, and the loss of diversity due to noise reduction is compensated by the up-sampling of the hard and diverse data points. We provide a case study in Appendix C.5 to further explain our argument.
@@ 4 Related Work
Pre-training is the critical phase for language models (LMs; 10, 60, 77, 15, 78, 79, 19) to obtain their foundation abilities for various downstream tasks. To improve pre-training, some works focus on data curation, such as adjusting domain mixing [86, 89], selecting valuable data points relevant to desired tasks [9, 20, 31], or transforming the instances based on downstream requirements [13, 28, 29]. Another line of work improves the optimization during pre-training by solving better data reweighting strategies [30], designing more effective optimizers [48, 69], or discovering better training recipes [39, 37]. Variations in model architectures [87] and training objectives [75] are also explored to boost pre-training stability and final LM performance. In this work, we utilize the knowledge from existing LMs to enhance pre-training.
@@
Given the high computational demands of large LMs during inference, there has been growing interest in pre-training small LMs [67]. However, achieving high performance with limited parameter sizes remains challenging because the training computation that small LMs need to match the capabilities of large LMs often exceeds Chinchilla-optimal [37] and scales as a power law with respect to the model size gap [40]. Despite these challenges, recent efforts have made promising progress by data quality improvement [52, 5, 91] or model pruning [58, 85]. We explore knowledge distillation as a complementary approach.
@@
Knowledge distillation (KD;36) uses a large teacher model to improve the performance of a small student model, which is widely used to build efficient neural network systems [62, 17, 64]. In NLP, early works primarily apply KD to encoder-only models [18, 49] for text classification by aligning token-level distribution [65], hidden states [72], and attention matrices [81, 82]. For generative LMs, a straightforward KD approach is training small LMs on the texts generated by large LMs [14, 63, 38]. Other works [32, 1, 83, 46, 84] explore better optimization objectives. However, these works focus on KD for fine-tuning LMs, while pre-training is critical for establishing core LM capabilities [2]. Therefore, we investigate KD in the pre-training stage to develop strong small base LMs.
@@ 5 Conclusion
One limitation of MiniPLM is that it requires the large LMs’ probabilities on texts from the pre-training corpus, making black-box KD for close-source LMs [60, 76] with MiniPLM challenging. For some APIs [59], this issue can be solved by specifying user-provided bias in the softmax operation, which allows obtaining the probabilities of given tokens one by one [12], at the expense of a large number of API calls.
@@
A promising future direction to explore is applying the difference-sampled corpus to pre-train LMs larger than the teacher LM, enabling weak-to-strong generalization [11]. Since data properties are critical for pre-training LMs across various sizes, the improvement of data diversity and difficulty is likely to be beneficial for LMs larger than the difference-sampling models.

### v3

@@ 3.1 Experimental Setup
We adopt the Qwen-1.5 (Bai et al., 2023) architecture in our experiments. We use the officially released 1.8B Qwen-1.5 model as the teacher LM and distill its knowledge into students with 200M, 500M, and 1.2B parameters. Detailed model configurations are provided in Appendix B.
@@
We construct pre-training corpora from the Pile (Gao et al., 2020). To control the computation in experiments, we pre-train all LMs on a maximum of 50B tokens, where documents are merged to construct instances with sequence lengths of 1,024. For online KD methods that incur additional train-time computation, we reduce their training steps to align the total training computation with pre-training without KD or offline KD methods. See Appendix B for more pre-training details.
@@
  • Vanilla KD (Muralidharan et al., 2024) minimizes the token-level forward KLD between $p$ and $q_{\bm{\theta}}$, which requires online inference of the teacher LM to obtain the token-level output distributions.
  • SeqKD (Kim & Rush, 2016) trains the student LM on the teacher-generated data. Since it is infeasible to generate all 50B tokens, we approximate Kim & Rush (2016) by using the first 768 tokens of each instance from the training corpus in Pre-Train w/o KD as the prompts and let the teacher LM generate the remaining tokens offline.
  • MiniLLM (Gu et al., 2024b) minimizes the reverse KLD between $p$ and $q_{\bm{\theta}}$ with PPO (Schulman et al., 2017), which requires online inference of the teacher LM and online sampling from the student LM. We treat the first 768 tokens of the instances from the training corpus of Pre-Train w/o KD as the prompts and sample 256 tokens from $q_{\bm{\theta}}$ during the exploration of PPO.
@@
We employ a 104M reference LM trained on 5B tokens. In Section 3.2 and 3.3, we consider a setting where $\mathcal{D}$ is sufficiently large, containing 105B tokens uniformly sampled from the Pile corpus. We reserve 5B tokens as $\mathcal{D}_{\text{ref}}$, and conduct Difference Sampling as per Eq. (4) on the other 100B tokens by setting $\alpha=0.5$ to construct a 50B-token corpus $\mathcal{D}^{\prime}$. In this way, the student LM is pre-trained on $\mathcal{D}^{\prime}$ for one epoch. We use the loss difference of the teacher and reference LM to sampled instances from $\mathcal{D}$, which is equivalent to Eq. (4) as all instances have 1,024 tokens. In Section 3.4, we evaluate MiniPLM in a data-limited setting, where $\mathcal{D}$ is controlled to contain 50B tokens and the student LM is trained on the difference-sampled data for multiple epochs. See Appendix D.3 for the ablation studies on the reference LM and sampling ratio $\alpha$.
@@
We assess the zero-shot accuracy of LMs trained with different methods on 9 downstream tasks widely used in examining the foundation abilities of base models (Touvron et al., 2023a; Groeneveld et al., 2024). We also test the language modeling capability of the LMs on a subset of DCLM (Li et al., 2024a), a high-quality corpus carefully curated with complex pipelines, to examine how well LMs capture broad and diverse knowledge. See Appendix B for more evaluation details.
@@ 1.8B Teacher $\rightarrow$ 500M Student
Table 1: Zero-shot accuracy scores on 9 widely-used downstream tasks and the average scores (Avg.). We use the Qwen-1.5 1.8B LM (Bai et al., 2023) as the teacher and Qwen LMs with 200M, 500M, and 1.2B parameters as the student. Student LMs with the same sizes consume the same training-time computation. The best scores of each model size are boldfaced.
@@ 3.2 Main Results
Figure 4: Language modeling loss on the DCLM (Li et al., 2024a) subset. We distill the knowledge of the 1.8B Qwen model (Bai et al., 2023) into student LMs from the Qwen family with 200M, 500M, and 1.2B parameters. We control the total training-time FLOPs of different methods to be the same.
@@
Table 2: Test loss predictions using the Scaling Law (Hoffmann et al., 2022). $N_{\text{stu}}$: the student LM size. $L_{\text{1T}}$, $L_{\text{10T}}$: the loss when Pre-Train w/o KD and MiniPLM process 1T and 10T tokens, with Vanilla KD consuming the same training FLOPs.
@@
Figure 4, compares the language modeling performance of the MiniPLM-trained LMs and baselines on the DCLM (Li et al., 2024a) subset, a diverse and high-quality dataset curated from web corpora. The results show that, given the same training-time FLOPs, LMs trained with MiniPLM achieve the lowest test losses. In Table 2, we extrapolate the test losses with the Scaling Law (Hoffmann et al., 2022) to simulate pre-training on 1T and 10T tokens (details in Appendix D.2), showing thatMiniPLM maintains its advantages at the scales of pre-training recent large LMs (Touvron et al., 2023a; Dubey et al., 2024). A critical stage in DCLM’s data-cleaning pipeline involves removing easy and common patterns, thereby enhancing challenging and diverse signals. Therefore, a lower test loss on DCLM suggests that, with the teacher LM’s guidance and the reference LM, the MiniPLM-trained LMs learn the diverse and hard knowledge better due to the up-sampling of the corresponding parts in the pre-training distribution, as illustrated in Figure 3(b).
@@ 3.3 KD Across Model Families
A noticeable advantage of MiniPLM over Vanilla KD and MiniLLM is its flexibility to distill the knowledge of a teacher LM into student LMs with completely different tokenizers and architectures without additional strategies like Boizard et al. (2024). In Table 5, we illustrate the performance when using the LMs from the Qwen (Bai et al., 2023) family as the teacher and the reference LM to distill knowledge into a 212M Llama3.1 (Dubey et al., 2024) model and a 140M Mamba (Gu & Dao, 2023) model. The results demonstrate the promising performance of MiniPLM in KD across model families, outperforming Pre-Train w/o KD and the existing offline KD baseline (SeqKD). This allows emerging LMs with novel architectures (Lieber et al., 2024; Sun et al., 2023) or advanced tokenization (Tay et al., 2022; Godey et al., 2022) to inherit knowledge from existing LMs, thereby facilitating the development of more efficient and higher-performed models.
@@ 3.4 Data-Limited Setting
We evaluate MiniPLM in a data-limited setting where $\mathcal{D}$ is constrained to contain 50B tokens. To this end, the student LM should be trained on the difference-sampled $\mathcal{D}^{\prime}$ over multiple epochs to ensure the total computation and trained tokens remain consistent with Pre-Train without KD. We split a $|\mathcal{D}_{\text{ref}}|$ containing 1B tokens to train the reference LM. Therefore, for a sampling ratio $\alpha$, the student LM should be trained for $\frac{|\mathcal{D}|}{\alpha|\mathcal{D}-\mathcal{D}_{\text{ref}}|}\approx\frac{1}{\alpha}$ epochs, given that $|\mathcal{D}_{\text{ref}}|\ll|\mathcal{D}|$. In Figure 5, we plot the loss curves of the 200M student LMs on the DCLM corpus when using $\alpha\in\left[0.5,0.25,0.125\right]$ and training the LM for around 2, 4, and 8 epochs, respectively. We can see that difference-sampling 25% data (with $\alpha=0.25$) and training the student LM for 4 epochs yields the best performance, which aligns with the observations in Muennighoff et al. (2023). The corpus sampled with a higher $\alpha$ does not achieve the best quality and diversity offered by Difference Sampling, while a lower $\alpha$ leads to rapid over-fitting of the student LM. These findings suggest that MiniPLM is a promising approach to enhance data utilization when high-quality web corpora become scarce (Villalobos et al., 2022). By extrapolating the loss curve of Pre-Train w/o KD using the Scaling Law in a data-constrained setting (Muennighoff et al., 2023), we estimate that it would require an additional 68B training tokens to match the performance of MiniPLM ($\alpha=0.25$, 4 epochs), which means MiniPLM reduces the pre-training data requirement by 2.4 times. See Appendix D.2 for more details on this extrapolation.
@@ 3.5 Analysis
Impact of Teacher Model.
Intuitively, larger teacher LMs will lead to better KD results, which is observed in recent works (Gu et al., 2024b). However, early works have also shown that an excessively large gap between teacher and student models can hinder effective KD (Mirzadeh et al., 2020). In Figure 6, we plot the performance of Vanilla KD and MiniPLM when distilling teacher LMs with different sizes into a 200M student model. We observe a similar phenomenon to Mirzadeh et al. (2020) on Vanilla KD and MiniPLM that larger teacher LMs are not necessarily more helpful for KD. However, we attribute this to different factors for Vanilla KD and MiniPLM. In Vanilla KD, the overhead introduced by larger LMs during training diminishes the benefits of distillation. As for MiniPLM, a 500M teacher LM proves most effective for distilling into a 200M student LM. Smaller teacher LMs (e.g., 300M) lack the capacity to identify the hard but valuable parts of the pre-training distribution, weakening the effectiveness of Difference Sampling as shown in Figure 3(b). On the other hand, the value scale of $\log p({\bm{x}})$ from oversized LMs (e.g., 4B) becomes too small compared to that of the reference LM, which tends to degenerate Difference Sampling into sampling with $p_{\text{ref}}({\bm{x}})$ only, losing the effect of the teacher LM. Future research could focus on optimizing teacher model size for MiniPLM or mitigating the impact of differing $\log p({\bm{x}})$ and $\log p_{\text{ref}}({\bm{x}})$ value scales.
@@
To further verify the effectiveness of Difference Sampling on improving the diversity of the pre-training corpus, in Table 6, we follow Friedman & Dieng (2023) to compute the semantic diversity of the difference-sampled corpus and the data used in other baseline pre-training approaches. The results show that the difference-sampled corpus has the highest diversity, despite that Difference Sampling is derived from minimizing the reverse KLD, which exhibits the mode-seeking behavior (Minka et al., 2005). We suspect the reason is that Difference Sampling down-samples the easy parts of the corpus containing repeated contents while up-sampling the hard parts consisting of diverse texts. These two components, with large $p({\bm{x}})$ values as seen in Figure 3(b), constitute the major modes of the teacher LM that $q_{\bm{\theta}}$ seeks during optimizing Eq. (1). Therefore, the mode-seeking behavior helps remove the noisy parts of the pre-training distribution, and the loss of diversity due to noise reduction is compensated by the up-sampling of the hard and diverse data points. We provide a case study in Appendix F to further explain our argument.
@@ 4 Related Work
Pre-training is the critical phase for language models (LMs; Brown et al., 2020; OpenAI, 2023; Team et al., 2024; Chowdhery et al., 2022; Touvron et al., 2023a; b; Dubey et al., 2024) to obtain their foundation abilities for various downstream tasks. To improve pre-training, some works focus on data curation, such as adjusting domain mixing (Xie et al., 2024; Ye et al., 2024), selecting valuable data points relevant to desired tasks (Brandfonbrener et al., 2024; Engstrom et al., 2024), or transforming the instances based on downstream requirements (Cheng et al., 2024; Gu et al., 2022; 2023). Another line of work improves the optimization during pre-training by solving better data reweighting strategies (Gu et al., 2024a), designing more effective optimizers (Liu et al., 2024; Shazeer & Stern, 2018), or discovering better training recipes (Hu et al., 2024; Hoffmann et al., 2022). Variations in model architectures (Xiong et al., 2020) and training objectives (Tay et al., 2023) are also explored to boost pre-training stability and final LM performance. In this work, we utilize the knowledge from existing LMs to enhance pre-training.
@@
Given the high computational demands of large LMs during inference, there has been growing interest in pre-training small LMs (Sardana et al., 2024). However, achieving high performance with limited parameter sizes remains challenging because the training computation that small LMs need to match the capabilities of large LMs often exceeds Chinchilla-optimal (Hoffmann et al., 2022) and scales as a power law with respect to the model size gap (Kaplan et al., 2020). Despite these challenges, recent efforts have made promising progress by data quality improvement (Mehta et al., 2024; Bellagente et al., 2024; Zhang et al., 2024) or model pruning (Muralidharan et al., 2024; Xia et al., 2024). We explore knowledge distillation as a complementary approach.
@@
Knowledge distillation (KD;Hinton et al., 2015) uses a large teacher model to improve the performance of a small student model, which is widely used to build efficient neural network systems (Park et al., 2019; Czarnecki et al., 2019; Salimans & Ho, 2022). In NLP, early works primarily apply KD to encoder-only models (Devlin et al., 2019; Liu et al., 2019) for text classification by aligning token-level distribution (Sanh et al., 2019), hidden states (Sun et al., 2019), and attention matrices (Wang et al., 2020; 2021). For generative LMs, a straightforward KD approach is training small LMs on the texts generated by large LMs (Chiang et al., 2023; Peng et al., 2023; Hsieh et al., 2023). Other works (Gu et al., 2024b; Agarwal et al., 2024; Wen et al., 2023; Li et al., 2024b; Wu et al., 2024) explore better optimization objectives. However, these works focus on KD for fine-tuning LMs, while pre-training is critical for establishing core LM capabilities (Allen-Zhu & Li, 2024). Therefore, we investigate KD in the pre-training stage to develop strong small base LMs.
@@ 5 Conclusion
One limitation of MiniPLM is that it requires the large LMs’ probabilities on texts from the pre-training corpus, making black-box KD for close-source LMs (OpenAI, 2023; Team et al., 2023) with MiniPLM challenging. For some APIs (OpenAI, 2022), this issue can be solved by specifying user-provided bias in the softmax operation, which allows obtaining the probabilities of given tokens one by one (Carlini et al., 2024), at the expense of a large number of API calls.
@@
A promising future direction to explore is applying the difference-sampled corpus to pre-train LMs larger than the teacher LM, enabling weak-to-strong generalization (Burns et al., 2023). Since data properties are critical for pre-training LMs across various sizes, the improvement of data diversity and difficulty is likely to be beneficial for LMs larger than the difference-sampling models.
Acknowledgements
This work is supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604) and Tsinghua University Initiative Scientific Research Program.